OpenStudy (anonymous):

If cos(theta) = 3/5, then sin(theta) can be what?

OpenStudy (anonymous):

\[\sin \theta=\frac{\sqrt{5^{2}-3^{2}}}{5}\]

OpenStudy (anonymous):

Using Pythagorean theorem to get the "opposite" side of the triangle. And opposite divided by hypotenuse is sine.

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

its a 3-4-5 right angled triangle

OpenStudy (anonymous):

\[\text{Sin}[\text{ArcCos}[3/5]]=\frac{4}{5} \]

OpenStudy (anonymous):

\[\cos^2 x+\sin^2 x=1\] so \[\sin^2 x=1-\cos^2 x\]

OpenStudy (anonymous):

you ok with these answers?

OpenStudy (anonymous):

Any valid method that comes up with 4/5 is fine and dandy.

OpenStudy (anonymous):

When I realized opposite/hypothenuse, I was hit with a long "OHHHHHHHHHH!"

OpenStudy (anonymous):

Rudy, how did you get your answer?

OpenStudy (anonymous):

How do you know that it is sqrt{5^2 - 3^2}

OpenStudy (anonymous):

That's the Pythagorean theorem. \[a^{2}=b^{2}+c^{2}\] Where a is the hypotenuse, b and c are the other two sides. Since you have the hypotenuse, just solve for b or c and you get: \[b^{2}=a^{2}-c^{2}\] and solving for b, you get: \[b=\sqrt{a^{2}-c^{2}}\]

OpenStudy (anonymous):

and 3,4 5 triangle 9these are lengths of the sides) is a right angled triangle because 5^2 = 3^2 + 4^2